# Search for the electromagnetic Dalitz decays $\chi_{cJ}\to e^{+}e^{-}\phi$

### v1

@@ 82 Zhengzhou University, Zhengzhou 450001, People’s Republic of China
a Deceased
b Also at the Moscow Institute of Physics and Technology, Moscow 141700, Russia
c Also at the Novosibirsk State University, Novosibirsk, 630090, Russia
d Also at the NRC ”Kurchatov Institute”, PNPI, 188300, Gatchina, Russia
e Also at Goethe University Frankfurt, 60323 Frankfurt am Main, Germany
f Also at Key Laboratory for Particle Physics, Astrophysics and Cosmology, Ministry of Education; Shanghai Key Laboratory for Particle Physics and Cosmology; Institute of Nuclear and Particle Physics, Shanghai 200240, People’s Republic of China
g Also at Key Laboratory of Nuclear Physics and Ion-beam Application (MOE) and Institute of Modern Physics, Fudan University, Shanghai 200443, People’s Republic of China
h Also at State Key Laboratory of Nuclear Physics and Technology, Peking University, Beijing 100871, People’s Republic of China
i Also at School of Physics and Electronics, Hunan University, Changsha 410082, China
j Also at Guangdong Provincial Key Laboratory of Nuclear Science, Institute of Quantum Matter, South China Normal University, Guangzhou 510006, China
k Also at MOE Frontiers Science Center for Rare Isotopes, Lanzhou University, Lanzhou 730000, People’s Republic of China
l Also at Lanzhou Center for Theoretical Physics, Lanzhou University, Lanzhou 730000, People’s Republic of China
m Also at the Department of Mathematical Sciences, IBA, Karachi 75270, Pakistan
n Also at Ecole Polytechnique Federale de Lausanne (EPFL), CH-1015 Lausanne, Switzerland
o Also at Helmholtz Institute Mainz, Staudinger Weg 18, D-55099 Mainz, Germany
p Also at Hangzhou Institute for Advanced Study, University of Chinese Academy of Sciences, Hangzhou 310024, China
@@ III DATA ANALYSIS
$\chi_{cJ}$ is from $\psi(3686)$ together with a photon, and the decay of interest is $\chi_{cJ}\to e^{+}e^{-}\phi$ with $\phi$ reconstructed from a charged kaon pair, so a signal event has 4 charged and 1 neutral tracks. Charged tracks detected in the MDC are required to be within a polar angle ($\theta$) range of $|\rm{cos\theta}|<0.93$, where $\theta$ is defined with respect to the $z$-axis, which is the symmetry axis of the MDC. The distance of closest approach to the interaction point (IP) must be less than 10 cm along the $z$-axis, and less than 1 cm in the transverse plane. The number of good charged tracks is required to be equal to four with zero net charge. Particle identification for charged tracks combines measurements of the energy deposited in the MDC ($dE/dx$) and the flight time in the TOF to form the probability values Prob($i$) ($i=e,\pi,K$). A track is considered as an electron when Prob(ee)>>Prob($\pi$) and Prob(ee)>>Prob($K$) and as a kaon when Prob(KK)>>Prob($\pi$) and Prob(KK)>>Prob($e$). Photon candidates are identified using isolated showers in the EMC. The deposited energy of each shower must be more than 25 MeV in the barrel region ($|\cos\theta|<0.80$) and more than 50 MeV in the end cap region ($0.86<|\cos\theta|<0.92$). To exclude showers that originate from charged tracks, the opening angle subtended by the EMC shower and the position of the closest charged track at the EMC must be greater than 10 degrees as measured from the IP. To suppress electronic noise and showers unrelated to the event, the difference between the EMC time and the event start time is required to be within [0, 700] ns. At least one photon is required in each event.
@@
The $\phi$ mass window is determined by fitting the $M_{K^{+}K^{-}}$ distribution of the signal MC sample with a double Gaussian function to describe the $\phi$ peak. The $\phi$ candidates are selected from $|M_{K^{+}K^{-}}-M_{\phi}|<8.5~\text{MeV}/c^{2}$, corresponding to about $2\sigma$. The $\chi_{cJ}$ mass windows are determined using the same method adopted for the $\phi$ mass window. The $\chi_{cJ}$ candidates are selected from $|M_{e^{+}e^{-}K^{+}K^{-}}-M_{\chi_{c0}}|<40.5~\text{MeV}/c^{2}$$|M_{e^{+}e^{-}K^{+}K^{-}}-M_{\chi_{c1}}|<17.5~\text{MeV}/c^{2}$$|M_{e^{+}e^{-}K^{+}K^{-}}-M_{\chi_{c2}}|<18.0~\text{MeV}/c^{2}$, respectively. The width of the $\chi_{cJ}$ mass windows corresponds to about $3\sigma$.
@@
The comparison of the remaining number of events in the signal region ($N_{obs}$), and the number of background events ($N_{bkg}$), estimated by the sideband method, is listed in Tab. 1.
Table 1: The comparison of $N_{obs}$ and $N_{bkg}$, where the uncertainties are estimated by assuming that the number of events follows a Poisson distribution.
@@ IV SYSTEMATIC UNCERTAINTIES
The signal MC events are generated without including the TFF correction. To determine the systematic uncertainties due to the TFFs, we generate the signal events using the TFF from the VMD, with the monopole approximation $F(q^{2})=1/(1-q^{2}/\Lambda^{2})$, with $\Lambda=3.773,~4.040~\rm{GeV}$ [15]. The maximum differences of signal efficiencies for the $\chi_{c0},~\chi_{c1}$, and $\chi_{c2}$ decays, which are 1.8%, 1.6%, and 1.1%, respectively, are taken as the corresponding systematic uncertainties caused by the TFFs.
[…]
@@ Acknowledgements.
The BESIII Collaboration thanks the staff of BEPCII (https://cstr.cn/31109.02.BEPC) and the IHEP computing center for their strong support. This work is supported in part by National Key R&D Program of China under Contracts Nos. 2023YFA1606000, 2023YFA1606704; National Natural Science Foundation of China (NSFC) under Contracts Nos. 11635010, 11935015, 11935016, 11935018, 12025502, 12035009, 12035013, 12061131003, 12192260, 12192261, 12192262, 12192263, 12192264, 12192265, 12221005, 12225509, 12235017, 12361141819, 12475093; the Chinese Academy of Sciences (CAS) Large-Scale Scientific Facility Program; CAS under Contract No. YSBR-101; 100 Talents Program of CAS; State Key Laboratory of Nuclear Physics and Technology, Peking University under Contracts No. NPT2025KFY03; The Institute of Nuclear and Particle Physics (INPAC) and Shanghai Key Laboratory for Particle Physics and Cosmology; German Research Foundation DFG under Contract No. FOR5327; Istituto Nazionale di Fisica Nucleare, Italy; Knut and Alice Wallenberg Foundation under Contracts Nos. 2021.0174, 2021.0299; Ministry of Development of Turkey under Contract No. DPT2006K-120470; National Research Foundation of Korea under Contract No. NRF-2022R1A2C1092335; National Science and Technology fund of Mongolia; National Science Research and Innovation Fund (NSRF) via the Program Management Unit for Human Resources & Institutional Development, Research and Innovation of Thailand under Contract No. B50G670107; Polish National Science Centre under Contract No. 2024/53/B/ST2/00975; Swedish Research Council under Contract No. 2019.04595; U. S. Department of Energy under Contract No. DE-FG02-05ER41374

### v2

@@ 82 Zhengzhou University, Zhengzhou 450001, People’s Republic of China
Deceased
a Also at the Moscow Institute of Physics and Technology, Moscow 141700, Russia
b Also at the Novosibirsk State University, Novosibirsk, 630090, Russia
c Also at the NRC ”Kurchatov Institute”, PNPI, 188300, Gatchina, Russia
d Also at Goethe University Frankfurt, 60323 Frankfurt am Main, Germany
e Also at Key Laboratory for Particle Physics, Astrophysics and Cosmology, Ministry of Education; Shanghai Key Laboratory for Particle Physics and Cosmology; Institute of Nuclear and Particle Physics, Shanghai 200240, People’s Republic of China
f Also at Key Laboratory of Nuclear Physics and Ion-beam Application (MOE) and Institute of Modern Physics, Fudan University, Shanghai 200443, People’s Republic of China
g Also at State Key Laboratory of Nuclear Physics and Technology, Peking University, Beijing 100871, People’s Republic of China
h Also at School of Physics and Electronics, Hunan University, Changsha 410082, China
i Also at Guangdong Provincial Key Laboratory of Nuclear Science, Institute of Quantum Matter, South China Normal University, Guangzhou 510006, China
j Also at MOE Frontiers Science Center for Rare Isotopes, Lanzhou University, Lanzhou 730000, People’s Republic of China
k Also at Lanzhou Center for Theoretical Physics, Lanzhou University, Lanzhou 730000, People’s Republic of China
l Also at the Department of Mathematical Sciences, IBA, Karachi 75270, Pakistan
m Also at Ecole Polytechnique Federale de Lausanne (EPFL), CH-1015 Lausanne, Switzerland
n Also at Helmholtz Institute Mainz, Staudinger Weg 18, D-55099 Mainz, Germany
o Also at Hangzhou Institute for Advanced Study, University of Chinese Academy of Sciences, Hangzhou 310024, China
@@ III DATA ANALYSIS
$\chi_{cJ}$ is from $\psi(3686)$ together with a photon, and the decay of interest is $\chi_{cJ}\to e^{+}e^{-}\phi$ with $\phi$ reconstructed from a charged kaon pair, so a signal event has 4 charged and 1 neutral tracks. Charged tracks detected in the MDC are required to be within a polar angle ($\theta$) range of $|\rm{cos\theta}|<0.93$, where $\theta$ is defined with respect to the $z$-axis, which is the symmetry axis of the MDC. The distance of closest approach to the interaction point (IP) must be less than 10 cm along the $z$-axis, and less than 1 cm in the transverse plane. The number of good charged tracks is required to be equal to four with zero net charge. Particle identification for charged tracks combines measurements of the energy deposited in the MDC ($dE/dx$) and the flight time in the TOF to form the probability values Prob($i$) ($i=e,\pi,K$). A track is considered as an electron when Prob(ee)>>Prob($\pi$) and Prob(ee)>>Prob($K$) and as a kaon when Prob(KK)>>Prob($\pi$) and Prob(KK)>>Prob($e$). The electron and kaon identification efficiencies are 90% and 94%, respectively. The probabilities of misidentifying $h$ as $h^{\prime}$, P($h\to h^{\prime}$), are estimated to be 0.2% for P($\pi\to e$), 0.2% for P($K\to e$), 4% for P($\pi\to K$) and 1% for P($e\to K$[44]. Photon candidates are identified using isolated showers in the EMC. The deposited energy of each shower must be more than 25 MeV in the barrel region ($|\cos\theta|<0.80$) and more than 50 MeV in the end cap region ($0.86<|\cos\theta|<0.92$). To exclude showers that originate from charged tracks, the opening angle subtended by the EMC shower and the position of the closest charged track at the EMC must be greater than 10 degrees as measured from the IP. To suppress electronic noise and showers unrelated to the event, the difference between the EMC time and the event start time is required to be within [0, 700] ns. At least one photon is required in each event.
@@
[…]
@@
The comparison of the remaining number of events in the signal region ($N_{\rm obs}$), and the number of background events ($N_{\rm bkg}$), estimated by the sideband method, is listed in Tab. 1.
Table 1: The comparison of $N_{\rm obs}$ and $N_{\rm bkg}$, where the uncertainties are estimated by assuming that the number of events follows a Poisson distribution.
@@ IV SYSTEMATIC UNCERTAINTIES
The signal MC events are generated with including the TFF correction. To determine the systematic uncertainties due to the TFFs, we generate the signal events using the TFF from the VMD, with the monopole approximation $F(q^{2})=1/(1-q^{2}/\Lambda^{2})$, with $\Lambda=3.773,~4.040~\rm{GeV}$ [15]. The maximum differences of signal efficiencies for the $\chi_{c0},~\chi_{c1}$, and $\chi_{c2}$ decays, which are 1.8%, 1.6%, and 1.1%, respectively, are taken as the corresponding systematic uncertainties caused by the TFFs.
[…]
@@ Acknowledgements.
The BESIII Collaboration thanks the staff of BEPCII (https://cstr.cn/31109.02.BEPC) and the IHEP computing center for their strong support. This work is supported in part by National Key R&D Program of China under Contracts Nos. 2025YFA1613900, 2023YFA1606000, 2023YFA1606704; National Natural Science Foundation of China (NSFC) under Contracts Nos. 11635010, 11935015, 11935016, 11935018, 12025502, 12035009, 12035013, 12061131003, 12192260, 12192261, 12192262, 12192263, 12192264, 12192265, 12221005, 12225509, 12235017, 12361141819, 12475093; the Chinese Academy of Sciences (CAS) Large-Scale Scientific Facility Program; CAS under Contract No. YSBR-101; 100 Talents Program of CAS; State Key Laboratory of Nuclear Physics and Technology, Peking University under Contracts No. NPT2025KFY03; The Institute of Nuclear and Particle Physics (INPAC) and Shanghai Key Laboratory for Particle Physics and Cosmology; German Research Foundation DFG under Contract No. FOR5327; Istituto Nazionale di Fisica Nucleare, Italy; Knut and Alice Wallenberg Foundation under Contracts Nos. 2021.0174, 2021.0299; Ministry of Development of Turkey under Contract No. DPT2006K-120470; National Research Foundation of Korea under Contract No. NRF-2022R1A2C1092335; National Science and Technology fund of Mongolia; National Science Research and Innovation Fund (NSRF) via the Program Management Unit for Human Resources & Institutional Development, Research and Innovation of Thailand under Contract No. B50G670107; Polish National Science Centre under Contract No. 2024/53/B/ST2/00975; Swedish Research Council under Contract No. 2019.04595; U. S. Department of Energy under Contract No. DE-FG02-05ER41374